1)You purchased 500 shares of stock for $28.50 a share. The initial margin requirement is 65 percent and the maintenance margin is 35 percent. What is the maximum percentage decrease that can occur in the stock price before you receive a margin call? Group of answer choices 38 percent 62 percent 57 percent 46 percent 35 percent

2)You recently purchased 200 shares of stock at a cost per share of $32.50. The initial margin requirement on this stock is 75 percent and the maintenance margin is 50 percent. The stock is currently valued at $35.00 a share. What is your current margin position? Ignore margin interest.

Group of answer choices

76.79 percent

75.69 percent

74.95 percent

73.83 percent

73.01 percent
Answer #1

1). P* = [{Shares Outstanding * Share Price * (1 - Initial Margin)} / Shares Outstanding] / (1 - Maintenance Margin)

= [{500 * $28.50 * (1 - 0.65)} / 500] / (1 - 0.35)

= [$4,987.50 / 500] / 0.65 = $9.975 / 0.65 = $15.3462

Maximum percentage decline = 1 - ($15.3462/$28.50) = 1 - 0.5385 = 0.4615, or 46 percent

Hence, 4th Option is correct.

2). Margin loan = Shares Outstanding * Share Cost * (1 - Initial Margin)

= 200 * $32.50 * (1 - 0.75) = $1,625

Current stock value = Shares Outstanding * Share Price = 200 * $35 = $7,000

Current equity = Current Stock Value - Margin Value = $7,000 - $1,625 = $5,375

Current margin = Current Equity / Current Stock Value = $5,375/$7,000 = 76.79 percent

Hence, 1st Option is correct.